Marigold Corp. thinks machine hours is the best activity base for its manufacturing overhead. The estimate of annual overhead costs for its jobs was $1550000. The company used 1000 hours of processing on Job No. B12 during the period and incurred overhead costs totaling $1600000. The budgeted machine hours for the year totaled 20000. How much overhead should be applied to Job No. B12?

Sheridan Company manufactures customized desks. The following pertains to Job No. 953:
Direct materials used $31800
Direct labor hours worked 600
Direct labor rate per hour $16.00
Machine hours used 500
Applied factory overhead rate per machine hour $30.00

What is the total manufacturing cost for Job No. 953?

$63900
$57800
$54800
$56400
$77500
$1600
$1550
Answer #1

1. Predetermined overhead rate = 1,550,000/20,000

= 77.50 per hour

Overhead applied = 77.50 * 1000

= 77,500

2. Total manufacturing overhead

= Direct Materials + Direct Labor + Overhead

= 31,800 + (600*16) + (500*30)

= 56,400